OpenStudy (lukecrayonz):

Use mathematical induction to prove the formula for every positive interger, n. 3+7+11+15+...+(4n-1)=n(2n+1)

OpenStudy (lukecrayonz):

@bahrom7893

OpenStudy (bahrom7893):

hmm do we prove that it's true for n-1 and n, or n and n+1?

OpenStudy (bahrom7893):

btw i have no clue ow to do this at the moment..

OpenStudy (bahrom7893):

@imranmeah91 take a look

OpenStudy (anonymous):

@myininaya

OpenStudy (bahrom7893):

lol myininaya will show up and say @amistre64 and then amistre will call for satellite.. lol just kiddin

OpenStudy (anonymous):

step 1 basecase 3 = 1(2 +1) yes assume true for some n prove true for n+1 3+7+11+15+...+(4n-1)=n(2n+1) consider the n+1 case add (4(n+1) -1) to each side = 4n +3 f(n+1) = n(2n+1) + 4n+3 = 2n^2 +n + 4n +3 =? (n+1)(2n +2 +1) =? (n+1) (2n+3) =? 2n^2 +5n +3 Yes

OpenStudy (lukecrayonz):

I got it :P

OpenStudy (lukecrayonz):

Kumar got it too!

myininaya (myininaya):

I was late :(